4g of sodium is reacted with excess oxygen to make sodium oxide. Work out the mass of sodium oxide produced.
1
Expert's answer
2020-10-26T14:48:40-0400

Solution:

The balanced chemical reaction:

4Na + O2 = 2Na2O

According to the chemical reaction: Moles of Na / 4 = Moles of Na2O / 2

n(Na)/2 = n(Na2O)


n(Na) = Mass of Na / Molar mass of Na

The molar mass of Na is 23 g mol-1.

n(Na) = 4g / 23 g mol-1 = 0.1739 mol


Hence,

n(Na2O) = n(Na)/2 = 0.1739 mol / 2 = 0.08695 mol.


n(Na2O) = Mass of Na2O / Molar mass of Na2O

The molar mass of Na2O is 62 g mol-1.

m(Na2O) = n(Na2O) × M(Na2O) = 0.08695 mol × 62 g mol-1 = 5.3909 g = 5.4 g

m(Na2O) = 5.4 g


Answer: 5.4 grams of sodium oxide (Na2O) produced.